# Preference-Based Gradient Estimation for ML-Based Approximate Combinatorial Optimization

### v1

@@ 5 Method
Figure 1 shows an overview of our approach. We augment an existing probabilistic approximation algorithm for a given CO problem using a GNN. The GNN receives the problem graph as input and produces a prior score for each edge. These scores are used as additional input alongside the graph for a parameterized version of an off-the-shelf CO approximation algorithm, which then produces a solution to the CO problem.
This same pipeline is used both during training and testing. Since the approximation algorithm is not differentiable in general, we use gradient estimation to obtain the gradients with respect to the GNN’s output. Existing gradient estimation schemes such as REINFORCE or I-MLE can be used for this, as introduced in subsection 3.3. We propose a new scheme for this purpose: preference-based gradient estimation (PBGE). The remainder of this section elaborates on our proposed training procedure as well as PBGE.
@@ 5.1 Parameterizing Approximation Algorithms
A given CO approximation algorithm that takes a problem graph as input can be parameterized by modifying the input graph, then running the approximation algorithm on the modified graph.
We calculate the modified graph ${\mathcal{G}}^{\prime}$ by using the GNN’s output to change the edge weights. Assume there is an arbitrary but fixed ordering of edges. The model outputs a prior score for each edge, $\bm{s}=f_{\bm{\theta}}({\mathcal{G}})\in\mathbb{R}^{|E|}$. A high score for a given edge is interpreted to mean that the respective edge should belong to the solution set with a higher probability mass. The approximation algorithms we use prefer including edges of low weight in the solution set. Therefore, we scale down the weights of edges that received high scores. Specifically, the edge weights are multiplied with $1-\sigma(\bm{s})$, where $\sigma$ is the element-wise sigmoid function. By running the CO approximation algorithm on this modified graph, we parameterize the approximation algorithm using the GNN’s output scores $\bm{s}$. In the formulas in the remainder of this paper, $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ represents the probability distribution defined by a probabilistic CO approximation algorithm parameterized in this way. It samples and outputs a vector $\hat{\bm{y}}\in\{0,1\}^{|E|}$ that represents a solution to the CO problem, such as a TSP tour or $k$-cut. A value of 1 in this vector indicates that the corresponding edge is in the solution set.
@@ 5.2 Preference-Based Gradient Estimation (PBGE)
In preference learning, a training instance consists of an input and a pair of possible outputs. The supervision signal is an annotation indicating that one of the outputs $\bm{y}_{w}$ is of higher quality than the other output $\bm{y}_{l}$. We can construct a similar setup for combinatorial optimization by leveraging a pre-existing probabilistic CO approximation algorithm $h$. Sampling from the approximation algorithm multiple times likely yields two solutions $\hat{\bm{y}}_{1},\hat{\bm{y}}_{2}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ of different quality for a given problem instance ${\mathcal{G}}\sim\mathcal{D}$ from dataset $\mathcal{D}$. These solutions can easily be ranked by applying the CO problem’s objective function [^2] Note that $J_{\mathrm{CO}}$ depends on a graph ${\mathcal{G}}$ and a predicted solution $\hat{\bm{y}}$, though we omit the graph parameter for readability $J_{\mathrm{CO}}$. This means assigning $\hat{\bm{y}}_{w}$ and $\hat{\bm{y}}_{l}$ such that $J_{\mathrm{CO}}(\hat{\bm{y}}_{w})<J_{\mathrm{CO}}(\hat{\bm{y}}_{l})$.
@@
$\begin{split}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\hat{\bm{y}}_{w},\hat{\bm{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim D}\\ &\left[d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})\log\left(\frac{h(\hat{\bm{y}}_{w}\mid{\mathcal{G}},\bm{s})}{h(\hat{\bm{y}}_{l}\mid{\mathcal{G}},\bm{s})}\right)\right].\end{split}$ (1)
Here, $d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})$ is a scaling factor [^3] The graph ${\mathcal{G}}$ is also a parameter to $d$, but we omit it in the notation for readability. As we will see later, its purpose is to scale the gradients based on the distance between the objective values of the losing and winning solutions.
Since we treat the CO approximation algorithm as a black box, we cannot calculate the probabilities $h(\hat{\bm{y}}_{w}\mid{\mathcal{G}},\bm{s})$ and $h(\hat{\bm{y}}_{l}\mid{\mathcal{G}},\bm{s})$ directly. We therefore introduce a proxy distribution $\pi(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})\approx h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ for which we can obtain probabilities directly. For all approximation algorithms we use in this paper, a high prior score in $\bm{s}$ for a certain edge increases the probability of this edge being included in the output $\hat{\bm{y}}$. This motivates the use of an exponential family distribution to model the proxy distribution $\pi$ for $h$:
$\pi(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})=\frac{\exp(-\langle\hat{\bm{y}},\bm{s}\rangle)}{\sum_{\bm{y}^{\prime}\in\mathcal{C}}\exp(-\langle\bm{y}^{\prime},\bm{s}\rangle)},$ (2)
@@
$\begin{split}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\bm{\hat{y}}_{w},\bm{\hat{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim D}\\ &\Big{[}d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})\big{(}\langle\bm{\hat{y}}_{l},\bm{s}\rangle-\langle\bm{\hat{y}}_{w},\bm{s}\rangle\big{)}\Big{]}.\end{split}$ (3)
Now, the gradient of this expectation w.r.t. $\bm{s}$ is
$\begin{split}\nabla_{\bm{s}}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\bm{\hat{y}}_{w},\bm{\hat{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim D}\\ &\big{[}d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})(\bm{\hat{y}}_{l}-\bm{\hat{y}}_{w})\big{]},\end{split}$ (4)
@@
$\begin{split}\nabla_{\bm{s}}\mathcal{L}({\mathcal{G}},\bm{s})&\approx d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})(\hat{\bm{y}}_{l}-\hat{\bm{y}}_{w}),\\ &\quad\text{where }\hat{\bm{y}}_{w},\hat{\bm{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}).\end{split}$ (5)
Intuitively, the gradient is $-1$ at a certain edge if that edge is in the better solution, but not in the worse solution. A gradient of $-1$ raises the GNN’s output score, meaning that the GNN will be nudged towards including this edge in its solution. Similarly, a gradient of $1$ means that the corresponding edge was in the worse solution, but not in the better solution. A positive gradient nudges the GNN’s output down, so it pushes the GNN towards not including this edge.
This gradient is similar to the ones used for preference learning with large language models (Rafailov et al., 2023; Meng et al., 2024): the gradient nudges the model to increase the likelihood of the better solution and to decrease the likelihood of the worse solution, and a scaling factor is used to weight important gradients more highly. Unlike the preference learning setting used with LLMs, we not only know which solution in a pair is better, but we can measure the quality of each solution exactly using the objective function. This means that we don’t need to rely on human annotators to rank pairs of examples. On top of this, we can leverage the objective function to more easily compute a suitable scaling factor.
@@
If the solutions $\hat{\bm{y}}_{w}$ and $\hat{\bm{y}}_{l}$ are of similar quality, we do not want to strongly nudge the GNN towards either solution. We therefore opt for scaling the gradient with the relative gap in quality between the two solutions,
@@
Note that this is always positive, because $J_{\text{CO}}(\hat{\bm{y}}_{l})>J_{\text{CO}}(\hat{\bm{y}}_{w})$. Using this scaling factor puts a larger emphasis on pairs of solutions where the difference in quality is large. In particular, if the two solutions are of the same quality, the gradient is set to zero, so we do not nudge the GNN towards either solution. The scaled gradient is
@@
The variance of the gradient can be decreased estimating the expectation in Equation 4 with more than one sample. However, this requires sampling two additional solutions from the approximation algorithm for each additional pair. We can increase the number of pairs more efficiently by first sampling a pool of solutions from the approximation algorithm, then constructing pairs from this pool. Since we want a gradient that nudges the model towards the best known solution, we form the pairs by combining the best solution from the pool with each of the weaker solutions.
In practice, the accuracy of the gradients depends heavily on the quality of the best found solution $\hat{\bm{y}}_{w}$. In our experiments, we noticed that early during training, the GNN cannot yet output good enough scores to consistently find reasonable $\hat{\bm{y}}_{w}$. To remedy this, we also run the approximation algorithm on the unmodified graph and add the resulting solutions to the pool from which the pairs are generated.
The complete training procedure is described in Algorithm 1.
@@
  $\bm{s}\leftarrow f_{\bm{\theta}}({\mathcal{G}})$// Evaluate the GNN
  // Sample $n$ solutions from approximation algorithm $h$, guided by GNN output $\bm{s}$, and $m$ solutions without guidance from $\bm{s}$
@@
  $\hat{\bm{y}}_{w}\leftarrow\operatorname*{arg\,min}_{\hat{\bm{y}}\in\{\hat{\bm{y}}_{1},\dots,\hat{\bm{y}}_{n+m}\}}J_{\mathrm{CO}}(\hat{\bm{y}})$
  $\nabla_{\bm{s}}\mathcal{L}\leftarrow\bm{0}$
  // Create pairs with $\hat{\bm{y}}_{w}$ and each other solution
  for $\hat{\bm{y}}_{l}\in\{\hat{\bm{y}}_{1},\dots,\hat{\bm{y}}_{n+m}\}$ do
     $\nabla_{\bm{s}}\mathcal{L}\leftarrow\nabla_{\bm{s}}\mathcal{L}+\left(\frac{J_{\text{CO}}(\hat{\bm{y}}_{l})}{J_{\text{CO}}(\hat{\bm{y}}_{w})}-1\right)\big{(}\hat{\bm{y}}_{l}-\hat{\bm{y}}_{w}\big{)}$
  end for
  Backpropagate gradient $\nabla_{\bm{s}}\mathcal{L}$
5.3 Decoding at Test Time
Regardless of how the model was trained, at test time, the model’s output needs to be converted (decoded) to a solution to the CO problem. This can simply be done by running the CO approximation algorithm with the model’s output as input, as described in section 4. The solution can be improved by running a probabilistic CO approximation algorithm repeatedly and using the best solution found as final output. Other decoders (like beam search) are also available, but haven’t shown promising results in preliminary experiments.
In the case of the Karger–Stein algorithm, we noticed empirically that simply modifying the input graph can lead to degenerate behavior during testing. The Karger–Stein algorithm uses the graph’s edge weights in two places: (1) when sampling an edge for contraction and (2) when comparing the cuts that resulted from different recursion arms. We noticed that the performance of our overall method can be improved when using a model trained with the setting described in subsection 5.2 by using the modified edge weights for the first case and the original edge weights for the second case. Intuitively, if the GNN makes a mistake when scaling the edge weights, using the original edge weights for comparing cuts can allow the Karger–Stein algorithm to find the optimal cut regardless.
For the algorithms we use in our experiments, we prove the following theoretical result, which shows that we can turn an approximation algorithm into an exact algorithm if we find an optimal modified input graph for Karger-Stein or random insertion.
Theorem 5.1.
Let $h$ be the Karger-Stein algorithm for the minimum $k$-cut problem or the random insertion algorithm for the TSP. Let $\bm{y}\in\{0,1\}^{|E|}$ be an optimal solution to the respective problem on graph ${\mathcal{G}}$. Then,

### v2

@@ 5 Method
We introduce a novel approach to CO by deriving a new gradient estimator that enables backpropagation through probabilistic approximation algorithms. This allows us, for the first time, to directly leverage solution quality rankings, which are automatically generated via objective function-based comparisons, to learn to improve CO approximation algorithms. We use graph neural networks (GNNs) to guide existing probabilistic approximation algorithms for a given CO problem. The GNN receives the problem graph as input and produces a prior score for each edge. These scores are used as additional input alongside the graph for a parameterized version of an off-the-shelf CO approximation algorithm, which then produces a solution to the CO problem. Figure 1 shows an overview of our approach.
Since the approximation algorithm is not differentiable in general, we use gradient estimation to obtain the gradients with respect to the GNN’s output. Existing gradient estimation schemes, such as REINFORCE, the straight-through estimator, Gumbel softmax, or I-MLE, either exhibit high bias or variance or require a differentiable loss function. We propose a new gradient estimation scheme based on preference-based optimization, which we term preference-based gradient estimation (PBGE).
@@ 5.1 Parameterizing approximation algorithms
An approximation algorithm that takes a problem graph as input can be indirectly parameterized by modifying the graph before executing the algorithm. The modified graph influences the behavior and output of the approximation process. We modify an input graph ${\mathcal{G}}^{\prime}$ by using the GNN’s output to change its edge weights. Assume there is an arbitrary but fixed ordering of edges. The model outputs for each edge a prior score $\bm{s}=f_{\bm{\theta}}({\mathcal{G}})\in\mathbb{R}^{|E|}$. A high score for a given edge is interpreted to mean that the respective edge should belong to the solution set with a higher probability mass. The approximation algorithms we use prefer including edges of low weight in the solution set. Therefore, we scale down the weights of edges that received high scores. Specifically, the edge weights are multiplied with $1-\sigma(\bm{s})$, where $\sigma$ is the element-wise sigmoid function. By running the CO approximation algorithm on this modified graph, we parameterize the approximation algorithm using the GNN’s output scores $\bm{s}$. In the remainder of this paper, $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})=h(\hat{\bm{y}}\mid{\mathcal{G}}^{\prime})$ denotes the probability distribution defined by a probabilistic CO approximation algorithm parameterized in this way. It samples and outputs a vector $\hat{\bm{y}}\in\{0,1\}^{|E|}$ that represents a solution to the CO problem, such as a TSP tour or $k$-cut. A value of $1$ in $\hat{\bm{y}}$ means that the corresponding edge is in the solution set.
@@ 5.2 Preference-based gradient estimation (PBGE)
In preference learning, a training instance consists of an input and a pair of possible outputs. The supervision signal is an annotation indicating that one of the outputs $\bm{y}_{w}$ is of higher quality than the other output $\bm{y}_{l}$. We can construct a similar setup for CO by leveraging a pre-existing probabilistic CO approximation algorithm $h$. Sampling from the approximation algorithm multiple times likely yields two solutions $\hat{\bm{y}}_{1},\hat{\bm{y}}_{2}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ of different quality for a given problem instance ${\mathcal{G}}\sim\mathcal{D}$ from dataset $\mathcal{D}$. These solutions can easily be ranked by applying the CO problem’s objective function  [^1] While $J_{\mathrm{CO}}$ depends on graph ${\mathcal{G}}$ and predicted solution $\hat{\bm{y}}$, we omit the graph parameter for readability. $J_{\mathrm{CO}}$. This means assigning $\hat{\bm{y}}_{w}$ and $\hat{\bm{y}}_{l}$ such that $J_{\mathrm{CO}}(\hat{\bm{y}}_{w})\leq J_{\mathrm{CO}}(\hat{\bm{y}}_{l})$.
@@
$\begin{split}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\hat{\bm{y}}_{w},\hat{\bm{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim\mathcal{D}}\left[d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})\log\left(\frac{h(\hat{\bm{y}}_{l}\mid{\mathcal{G}},\bm{s})}{h(\hat{\bm{y}}_{w}\mid{\mathcal{G}},\bm{s})}\right)\right].\end{split}$ (1)
Here, $d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})$ is a scaling factor. [^2] The input graph ${\mathcal{G}}$ is also an argument of $d$, but we omit it for ease of readability. As we will see later, its purpose is to scale the gradients based on the distance between the objective values of $\hat{\bm{y}}_{w}$ and $\hat{\bm{y}}_{l}$.
Since we treat the CO approximation algorithm as a black box, we cannot calculate the probabilities $h(\hat{\bm{y}}_{w}\mid{\mathcal{G}},\bm{s})$ and $h(\hat{\bm{y}}_{l}\mid{\mathcal{G}},\bm{s})$ directly. We therefore introduce a proxy distribution $\pi(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})\approx h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ for which we can obtain probabilities directly. For all approximation algorithms used in this paper, a high prior score in $\bm{s}$ for a certain edge increases the probability of this edge being included in the output $\hat{\bm{y}}$. This motivates the use of an exponential family distribution to model the proxy distribution $\pi$ for $h$:
$\pi(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})=\frac{\exp(\langle\hat{\bm{y}},\bm{s}\rangle)}{\sum_{\bm{y}^{\prime}\in\mathcal{C}}\exp(\langle\bm{y}^{\prime},\bm{s}\rangle)},$ (2)
@@
$\begin{split}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\bm{\hat{y}}_{w},\bm{\hat{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim D}\Big{[}d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})\big{(}\langle\bm{\hat{y}}_{l},\bm{s}\rangle-\langle\bm{\hat{y}}_{w},\bm{s}\rangle\big{)}\Big{]}.\end{split}$ (3)
Now, the gradient of this expectation with respect to $\bm{s}$ is
$\begin{split}\nabla_{\bm{s}}\mathcal{L}(\mathcal{D},\bm{s})=&\;\mathbb{E}_{\bm{\hat{y}}_{w},\bm{\hat{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}),\>{\mathcal{G}}\sim D}\big{[}d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})(\bm{\hat{y}}_{l}-\bm{\hat{y}}_{w})\big{]},\end{split}$ (4)
@@
$\begin{split}\nabla_{\bm{s}}\mathcal{L}({\mathcal{G}},\bm{s})&\approx d(\hat{\bm{y}}_{w},\hat{\bm{y}}_{l})(\hat{\bm{y}}_{l}-\hat{\bm{y}}_{w}),\quad\text{where }\hat{\bm{y}}_{w},\hat{\bm{y}}_{l}\sim h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s}).\end{split}$ (5)
Intuitively, the gradient is negative at a certain edge if that edge is in the better solution, but not in the worse solution. A negative gradient raises the GNN’s output score, meaning that the GNN will be nudged towards including this edge in its solution. Similarly, a positive gradient means that the corresponding edge was in the worse solution, but not in the better solution. This nudges the GNN’s output down, so it pushes the GNN towards not including this edge.
The form of the gradient is reminiscent of those used for preference learning with large language models (Rafailov et al., 2023; Meng et al., 2024). There, the gradient makes the model increase the likelihood of the better solution and decrease the likelihood of the worse solution, and a scaling factor is used to weight important gradients more highly. Unlike the preference learning setting used with LLMs, we not only know which solution in a pair is better, but we can measure the quality of each solution exactly using the objective function. This eliminates the need for human annotators to rank pairs of examples. Moreover, we can leverage the objective function to more easily compute a suitable scaling factor.
@@
If the solutions $\hat{\bm{y}}_{w}$ and $\hat{\bm{y}}_{l}$ are of similar quality, we do not want to strongly move the GNN towards either solution. We therefore scale the gradient with the relative optimality gap between the two solutions,
@@
This is always non-negative, since, by definition, $J_{\text{CO}}(\hat{\bm{y}}_{l})\geq J_{\text{CO}}(\hat{\bm{y}}_{w})$. Using this scaling factor places more weight on pairs of solutions where the difference in their respective optimality gaps is large. In particular, if the two solutions are of the same quality, the gradient is set to zero, so we do not move the GNN towards either solution. The scaled gradient is
@@
$\bm{s}\leftarrow f_{\bm{\theta}}({\mathcal{G}})$
$\triangleright$ Sample $n$ solutions from $h$ guided by $\bm{s}$, and $m$ unguided solutions
@@
$\hat{\mathcal{Y}}\leftarrow\{\hat{\bm{y}}_{1},\dots,\hat{\bm{y}}_{n+m}\}$
$\hat{\bm{y}}_{w}\leftarrow\operatorname*{arg\,min}_{\hat{\bm{y}}\in\hat{\mathcal{Y}}}J_{\mathrm{CO}}(\hat{\bm{y}})$
$\triangleright$ Estimate gradient using each pair involving $\hat{\bm{y}}_{w}$
$\nabla_{\bm{s}}\mathcal{L}({\mathcal{G}},\bm{s})\approx\displaystyle\sum_{\hat{\bm{y}}_{l}\in\hat{\mathcal{Y}}}\!\left(\frac{J_{\text{CO}}(\hat{\bm{y}}_{l})}{J_{\text{CO}}(\hat{\bm{y}}_{w})}-1\right)\!\big{(}\hat{\bm{y}}_{l}-\hat{\bm{y}}_{w}\big{)}$
Backpropagate gradient $\nabla_{\bm{s}}\mathcal{L}({\mathcal{G}},\bm{s})$
The variance of the gradient can be reduced by estimating the expectation in Equation 4 by creating a pool of solutions with the approximation algorithm and by constructing pairs from this pool. We form the pairs by combining the best solution from the pool with each of the weaker solutions. In practice, the accuracy of the gradients depends heavily on the quality of the best found solution $\hat{\bm{y}}_{w}$. At the beginning of training, the GNN cannot yet output good enough scores to consistently find reasonable $\hat{\bm{y}}_{w}$. To remedy this, we also run the approximation algorithm on the unmodified graph and add the resulting solutions to the pool from which the pairs are generated. The complete training procedure is described in Algorithm 1.
At test time, the model’s output needs to be converted (decoded) to a solution to the CO problem. This can simply be done by running the CO approximation algorithm with the model’s output as input, as described in Section 4. The solution can be improved by running a probabilistic CO approximation algorithm repeatedly and using the best solution found as final output.
5.3 Theoretical analysis
We prove the following theoretical results, which characterize under what conditions we can turn an approximation algorithm into an exact algorithm if we find an optimal modified input graph. First, we prove this for the Karger–Stein algorithm.
Theorem 1.
Let $h(\hat{\bm{y}}\mid{\mathcal{G}},\bm{s})$ be the Karger–Stein algorithm guided by scores $\bm{s}\in\mathbb{R}^{|E|}$, and let $\bm{y}\in\{0,1\}^{|E|}$ be a minimum $k$-cut on graph ${\mathcal{G}}=(V,E)$. Then, $\lim_{\sigma(\bm{s})\to\bm{y}}h(\bm{y}\mid{\mathcal{G}},\bm{s})=1.$
Now, we consider an arbitrary CO problem whose solutions are a set of edges of minimum total weight that satisfy the constraints (e.g.,  minimum $k$-cut or TSP).
Theorem 2.
Let $J_{\text{CO}}$ be the objective of a minimization problem on graphs with $J_{\text{CO}}(\hat{\bm{y}})=\sum_{e\in\hat{\bm{y}}}w(e)$ and $w(e)>0$ and let $\bm{y}\in\{0,1\}^{|E|}$ be an optimal solution on graph ${\mathcal{G}}$. Moreover, let $h(\hat{\bm{y}}\mid{\mathcal{G}})$ be a probabilistic approximation algorithm with a bound on the relative error, that is, there exists a function $\alpha$ such that for all $\hat{\bm{y}}\sim h(\hat{\bm{y}}\mid{\mathcal{G}})$, $J_{\text{CO}}(\hat{\bm{y}})\!\leq\!\alpha({\mathcal{G}})J_{\text{CO}}(\bm{y})$. Then, $\lim_{\sigma(\bm{s})\to\bm{y}}h(\bm{y}\mid{\mathcal{G}},\bm{s})=1.$
In other words, in the limit, the approximation algorithm guided by $\bm{s}$ always finds optimal solution $\bm{y}$. From this, the following result regarding insertion algorithms on TSP immediately follows:
Corollary 3.
Let ${\mathcal{G}}=(V,E,w)$ be an undirected graph, and let $h(\hat{\bm{y}}\mid{\mathcal{G}})$ be a probabilistic insertion algorithm, such as random insertion. Let $\bm{y}\in\{0,1\}^{|E|}$ be a minimal length TSP tour on ${\mathcal{G}}$. Then,